OpenStudy (anonymous):

2 square root of 7x + square root of 7x - 3 square root of 7x - 7

OpenStudy (anonymous):

is it ... \[\LARGE 2\sqrt{7x}+\sqrt{7x}-3 +\sqrt{7x}-7\] or \[\LARGE 2\sqrt{7x}+\sqrt{7x-3} +\sqrt{7x-7}\] ?

OpenStudy (anonymous):

yes to the first one

OpenStudy (anonymous):

then I think this should help... \[\Large a\sqrt{7x}+b\sqrt{7x}=(a+b)\sqrt{7x}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

do you know how to do it now ?

OpenStudy (anonymous):

i am trying to figure it out right now

OpenStudy (anonymous):

If you can't just let me know... I'll help you ;)

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

you're welcome ;)

OpenStudy (anonymous):

is the answer 3 square root of 7x - 7

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\Large 2\sqrt{7x}+\sqrt{7x}-3+\sqrt{7x}-7=\] \[\Large 2\sqrt{7x}+\sqrt{7x}+\sqrt{7x}-7-3=\] \[\Large (2+1+1)\sqrt{7x}-7-3=\] \[\Large 4\sqrt{7x}-10\] @cassandre if you're still confused about steps I made solving it , feel free to ask ;)

OpenStudy (anonymous):

omg, i see where the issue is, i forgot to combine the -7 with the - 3

OpenStudy (anonymous):

thanks for helping

OpenStudy (anonymous):

It happens don't worry ;) .. Glad to help :)